OpenStudy (anonymous):

Temperature measurements are converted between degrees Fahrenheit (F) and degrees Celsius (C) by the formula F = (9/5)C + 32°. The temperature in Dallas rose from the morning temperature of 49°F to the afternoon high of 75°F. What was the difference in temperature in degrees Celsius, to the nearest degree? Im thinking 26 but seems too simple to solve

OpenStudy (anonymous):

\[\frac59 (F-32)=C\]

OpenStudy (anonymous):

you need to convert from F to C first

OpenStudy (campbell_st):

rewrite the equation to make C the subject \[C = 5/9(F - 32)\] now do the substitutions

OpenStudy (anonymous):

C = (F -32)5/9 75-49 = 26 F = (26-32)5/9 C

OpenStudy (anonymous):

what cinar said \[\frac59 (49-32)=C\] and \[\frac59 (75-32)=C\] and compare those two

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

i got 14

OpenStudy (anonymous):

no yes?